# Split Two-Periodic Aztec Diamond

### v4

@@ 1.1. The Aztec Diamond and Dimer Models
Over the past quarter century, planar dimer models have been an active area of study. A dimer covering of a graph, is a subset of the edges such that each vertex of the graph is incident to exactly one edge in the covering. A dimer model is a probability measure on the set of all dimer coverings. To define the measure one can assign edge weights to the graph. The probability of a certain covering is proportional to the product of the edge weights of the dimers contained in the covering. The ideas behind dimer models were initially introduced by Kasteleyn [16] and Temperley and Fisher [20].
Dimer models are often studied under some appropriate scaling limit. Cohn, Kenyon and Propp [10] showed that a class of uniform dimer models satisfy a variational principle and that the height function of the model converges, in the scaling limit, to a deterministic limit. Subsequently, Kenyon, Okounkov, and Sheffield [18] showed that there are three types of Gibbs measures that can appear in dimer model with doubly periodic edge weights. These macroscopic regions are classified as: frozen, rough, and smooth. [^1] In the literature these regions are also referred to as solid, liquid, and gas, respectively. In the frozen region dimers are deterministic. In the rough region dimer correlations decay polynomially with distance, while in the smooth region dimer correlations decay exponentially with distance. Not all dimer models exhibit smooth regions, however most exhibit rough regions.
A seminal example of a dimer model is domino tilings of the Aztec diamond [12]. The asymptotics of the simplest case, the Aztec diamond with uniform edge weights, was studied by Jockusch, Propp and Shor in [14]. They proved that the boundary between the deterministic and non-deterministic regions forms a circle. This is known as the Arctic Circle Theorem. A key aspect to understanding correlations of the Aztec diamond is the fact that the model is a determinantal point process. One approach to computing correlations is through computing the inverse Kasteleyn matrix. [^2] For a general introduction to dimer models via Kasteleyn theory see [17]. The inverse Kasteleyn matrix for the Aztec diamond with two-periodic weights was originally computed by Chhita and Young in [9]. Further asymptotics of this model have been studied in [8, 7, 1] among others. The two-periodic Aztec diamond is, notably, the simplest Aztec diamond model which exhibits all three macroscopic regions.
@@
An alternative approach to understanding the correlations of the dimer model uses a bijection between the Aztec diamond and a non-intersecting paths model along side the Eynard-Mehta theorem [13]. The correlation kernel of the two-periodic Aztec diamond was computed using this approach by Duits and Kuijlaars [11], via matrix valued orthogonal polynomials and Berggren and Duits [3], via a Wiener-Hopf factorization. In both of these instances, the methods produced forms of the correlation kernel that are well suited for asymptotic analysis. The Wiener-Hopf method has been related to the Kasteleyn treatment in [6] and to the MVOP method in [19].
Recently, there have been many results about the Aztec diamond with more general periodic weightings, see [5, 4, 2]. Notably in [5], Boutillier and de Tilière compute the inverse Kasteleyn matrix for the Aztec diamond with Fock’s weights, which are gauge equivalent to any positive weighting of the model, and study the limit shape in cases where the weights are periodic. There is less literature, however, regarding the asymptotics of weightings that are not periodic in nature, which is the main goal of this work.
In Section 1.2, we define an extension of the typical two-periodic Aztec diamond, which we refer to as the split two-periodic Aztec diamond. In the split two-periodic model, periodicity is broken along a line which we refer to as the interface of the model. A simulation of the split two-periodic Aztec diamond is shown in Figure 2. We compute the correlation kernel of this new model by extending the methods developed in [3]. The statement of the correlation kernel is found in Section 2.1. The rest of our discussion in Section 2 relates to defining the macroscopic regions and computing the local asymptotic of the model away from the boundaries and interface.
Figure 2. Simulated split Aztec diamond where $N=200$, $\alpha=1/4$, and $\beta=1/2$. The tiles are colored by an 8 color gray scale to accentuate the smooth region. Original code for simulation was provided by Sunil Chhita.
@@ 1.2. The Split Two-Periodic Aztec Diamond
(2) $\mathtt{B}^{\text{Az}}_{n}=\Big{\{}(2j,2k+1):0\leq j\leq n\text{ and }0\leq k\leq n-1\Big{\}}$
The Aztec diamond is a bipartite graph, so we write the set of all vertices as $\mathtt{V}_{n}^{\text{Az}}=\mathtt{W}^{\text{Az}}_{n}\oplus\mathtt{B}^{\text{Az}}_{n}$. The edges are given by,
(3) $\mathtt{E}^{\text{Az}}_{n}=\Big{\{}((2j+1,2k),(2j+1\pm 1,2k+1)):0\leq j\leq n-1,0\leq k\leq n-1\Big{\}}\\ \cup\Big{\{}((2j+1,2k),(2j+1\pm 1,2k-1)):0\leq j\leq n-1,1\leq k\leq n\Big{\}}$
0000111111112222222233333333444444445555555566666666
00001111111122222222333333334444444455555555666666667777777788888888
@@
(4) $\text{wt}(b,w)=\begin{cases}\alpha^{2}&\text{if }w_{x}\equiv 1\bmod{4},\,w_{y}\equiv 0\bmod{4},\,w-b=(\pm 1,-1),\text{ and }w_{x}<2N\\ \alpha^{-2}&\text{if }w_{x}\equiv 1\bmod{4},\,w_{y}\equiv 2\bmod{4},\,w-b=(\pm 1,-1),\text{ and }w_{x}<2N\\ \beta^{2}&\text{if }w_{x}\equiv 1\bmod{4},\,w_{y}\equiv 0\bmod{4},\,w-b=(\pm 1,-1),\text{ and }w_{x}>2N\\ \beta^{-2}&\text{if }w_{x}\equiv 1\bmod{4},\,w_{y}\equiv 2\bmod{4},\,w-b=(\pm 1,-1),\text{ and }w_{x}>2N\\ 1&\text{if }w_{x}\equiv 1\bmod{4},\text{ and }w-b=(\pm 1,1)\\ 1&\text{if }w_{x}\equiv 3\bmod{4},\text{ and }w-b=(\pm 1,\pm 1)\\ 0&\text{otherwise}\end{cases}$
where $10<\alpha,\beta<$ 1 are constants. The last case in the above formula indicates the the weight is 0 if the vertices are not neighbors. If $\alpha=\beta$, this model recovers the usual two-periodic Aztec diamond. If $\alpha=\beta=1$ we recover the uniform Aztec diamond, however we will generally assume that the constants are not equal to 1. The split two-periodic Aztec diamond of size $42N=$ 4 is shown in Figure 4. We will refer to the line were the weighting changes as the interface of the model. In this coordinate convention, the interface occurs along the vertical line $x=2N$.
@@
Figure 4. The Aztec diamond of size $42N=$ 4 with coordinates given by equations (1)-(2) and weights given by equation (4). All unlabeled edged have weight 1.
To produce the correlation kernel for the split two-periodic Aztec diamond we will extend the work of Berggren and Duits in [3]. Their work computes the correlation kernel for a class of non-intersecting paths models, which are detailed in Section 5 of this work. Their work relates the Aztec diamond to a non-intersecting paths process, which is shown to be equivalent to the Aztec diamond model. We detail this equivalence in Section 5 of this work. For now, we make some simple changes to the coordinate convention on the Aztec diamond which will allow us to immediately relate the statement of the kernel to exact positions in the Aztec diamond. For details on why this transformation is valid, one should refer to the bijections developed in Section 5.
To transform the coordinates, we start by moving the white vertices of the Aztec diamond to be level with the row of black vertices immediately below them. Consider $w\in\mathtt{W}$ with initial coordinate $(2j+1,2k)$, it now has the coordinate $(2j+1,2k-1)$. Next, to any vertex $v=(v_{x},v_{y})$ we make the following change to the $y$-coordinate,
$v_{y}\mapsto\frac{1}{2}v_{y}-2N-\frac{1}{2}$
This transformation is depicted in Figure 5. For the Aztec diamond of size $n$ the vertex sets become the following under the the new coordinate convention,
(5) $\mathtt{W}^{\text{Az-Alt}}_{n}=\Big{\{}(2j+1,k):0\leq j<n\text{ and }-n-1\leq k\leq-1\Big{\}}$
(6) $\mathtt{B}^{\text{Az-Alt}}_{n}=\Big{\{}(2j,k):0\leq j<n\text{ and }-n\leq k\leq-1\Big{\}}$
And the edges are now given by,
(7) $\mathtt{E}^{\text{Az-Alt}}_{n}=\Big{\{}\left((2j+1,k),(2j+1\pm 1,k)\right):0\leq j<n\text{ and }-n\leq k\leq-1\Big{\}}\\ \cup\Big{\{}\left((2j+1,k),(2j+1\pm 1,k+1)\right):0\leq j<n\text{ and }-n-1\leq k\leq-2\Big{\}}$
For the correlation kernel given in Section 2.1, these transformed coordinates will be the ones we use.
0000111111112222222233333333444444445555555566666666
0011112222333344445555666611-1- 122-2- 233-3- 344-4- 4
Figure 5. The above depicts the coordinate changes on the size $3$ Aztec diamond.
@@ 2. Summary of Results
In this section, we will detail the main results of this work. This starts with the statement of the correlation kernel of the split two-periodic Aztec diamond, which we later compute by extending the methods of Berggren and Duits [3]. We then relate this kernel to the kernel of the typical two-periodic Aztec diamond model. Lastly, we classify the macroscopic regions of the model by way of computing the local asymptotics. In particular, we emphasize the differences in the local asymptotics of the split two-periodic model compared to the typical two-periodic model.
@@ 2.1. An Integral Representation of the Correlation Kernel
$\phi_{\varepsilon,1}(z)$ $=\begin{pmatrix}1&\varepsilon^{2}z^{-1}\\ \varepsilon^{-2}&1\end{pmatrix}$ $\phi_{\varepsilon,2}(z)$ $=\frac{1}{1-z^{-1}}\begin{pmatrix}1&\varepsilon^{2}z^{-1}\\ \varepsilon^{-2}&1\end{pmatrix}$
$\phi_{3}(z)$ $=\begin{pmatrix}1&z^{-1}\\ 1&1\end{pmatrix}$ $\phi_{4}(z)$ $=\frac{1}{1-z^{-1}}\begin{pmatrix}1&z^{-1}\\ 1&1\end{pmatrix}$
@@
(8) $\Phi_{\varepsilon}(z)=\phi_{\varepsilon,1}(z)\phi_{\varepsilon,2}(z)\phi_{3}(z)\phi_{4}(z)$
@@
(9) $\Phi_{\varepsilon}(z)=E_{\varepsilon}(z)\begin{pmatrix}r_{\varepsilon,1}(z)&0\\ 0&r_{\varepsilon,2}(z)\end{pmatrix}E_{\varepsilon}(z)^{-1}$
@@
(11) $r_{\varepsilon,2}(z)=\frac{1}{(z-1)^{2}}\left((z+1)^{2}+2z(\varepsilon^{2}+\varepsilon^{-2})-2(\varepsilon+\varepsilon^{-1})\sqrt{z^{3}+(\varepsilon^{2}+\varepsilon^{-2})z^{2}+z}\right)$
We also adopt the notation,
@@
$=\begin{pmatrix}\frac{1}{2}+\frac{z(\varepsilon^{2}-1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}&\frac{\varepsilon^{2}(z+1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}\\ \frac{z(z+1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}&\frac{1}{2}-\frac{z(\varepsilon^{2}-1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}\end{pmatrix}$
Lastly we define the function,
(14) $g_{\alpha,\beta}(z)=\frac{2z(1+\alpha^{2}\beta^{2})+\beta^{2}(z^{2}+1)+\alpha^{2}(z^{2}+1)}{4\sqrt{(z+\alpha^{2})(1+\alpha^{2}z)(z+\beta^{2})(1+\beta^{2}z)}}$
Note that $g_{\alpha,\beta}(z)=g_{\beta,\alpha}(z)$. We are now prepared to state the main theorem of this paper,
@@
Let $-N\leq\xi,\xi^{\prime}\leq-1$ and $0<m<N$. The split two-periodic Aztec diamond of size $2N$ has correlation kernel given by
(15) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\alpha}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\$
@@
(16) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(w)F_{\alpha,2}(w)F_{\beta,2}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
when $\nicefrac{{N}}{{2}}<m^{\prime}<N$. For both cases, if $m\leq\nicefrac{{N}}{{2}}$ then $\varepsilon=\alpha$ and if $m>\nicefrac{{N}}{{2}}$ then $\varepsilon=\beta$. Additionally, $\gamma_{1}$ is a contour surrounding $1$ and $\gamma_{0,1}$ is a contour surrounding 00 and $\gamma_{1}$. Both contours are oriented in the positive direction.
$\gamma_{1}$$\gamma_{0,1}$

### v5

@@ 1.1. The Aztec Diamond and Dimer Models
Over the past quarter century, planar dimer models have been an active area of study. A dimer covering of a graph, is a subset of the edges such that each vertex of the graph is incident to exactly one edge in the covering. A dimer model is a probability measure on the set of all dimer coverings. To define the measure one can assign edge weights to the graph. The probability of a certain covering is proportional to the product of the edge weights of the dimers contained in the covering. The ideas behind dimer models were initially introduced by Kasteleyn [Kas61] and Temperley and Fisher [TF61].
Dimer models are often studied under some appropriate scaling limit. Cohn, Kenyon and Propp [CKP01] showed that a class of uniform dimer models satisfy a variational principle and that the height function of the model converges, in the scaling limit, to a deterministic limit. Subsequently, Kenyon, Okounkov, and Sheffield [KOS03] showed that there are three types of Gibbs measures that can appear in dimer model with doubly periodic edge weights. These macroscopic regions are classified as: frozen, rough, and smooth. [^1] In the literature these regions are also referred to as solid, liquid, and gas, respectively. In the frozen region dimers are deterministic. In the rough region dimer correlations decay polynomially with distance, while in the smooth region dimer correlations decay exponentially with distance. Not all dimer models exhibit smooth regions, however most exhibit rough regions.
A seminal example of a dimer model is domino tilings of the Aztec diamond [EKLP92]. The asymptotics of the simplest case, the Aztec diamond with uniform edge weights, was studied by Jockusch, Propp and Shor in [JPS98]. They proved that the boundary between the deterministic and non-deterministic regions forms a circle. This is known as the Arctic Circle Theorem. A key aspect to understanding correlations of the Aztec diamond is the fact that the model is a determinantal point process. One approach to computing correlations is through computing the inverse Kasteleyn matrix. [^2] For a general introduction to dimer models via Kasteleyn theory see [Ken2009lectures]. The inverse Kasteleyn matrix for the Aztec diamond with two-periodic weights was originally computed by Chhita and Young in [CY14]. Further asymptotics of this model have been studied in [CJY15, CJ16, BCJ18] among others. The two-periodic Aztec diamond is, notably, the simplest Aztec diamond model which exhibits all three macroscopic regions.
@@
An alternative approach to understanding the correlations of the dimer model uses a bijection between the Aztec diamond and a non-intersecting paths model along side the Eynard-Mehta theorem [EM97]. The correlation kernel of the two-periodic Aztec diamond was computed using this approach by Duits and Kuijlaars [DK21], via matrix valued orthogonal polynomials (MVOP) and Berggren and Duits [BD19], via a Wiener-Hopf factorization. In both of these instances, the methods produced forms of the correlation kernel that are well suited for asymptotic analysis. The Wiener-Hopf method has been related to the Kasteleyn treatment in [CD23] and to the MVOP method in [KP24].
Recently, there have been many results about the Aztec diamond with more general periodic weightings, see [BB24, Ber19, BD23, BB23]. There is less work, however, on dimer models that involve non-periodic weightings. One notable non-periodic set up is the $q^{vol}$-weightings [CY14, BBCCR17]. Additionally in [BdT24] Boutillier and de Tilière compute the inverse Kasteleyn matrix for the Aztec diamond with Fock’s weights, which are only quasi-periodic. Other set ups and approaches, such as $t$-embeddings [CLR23, BNR24] and the Enyard-Mehta theorem [DK21, BD19], seem to have possible applications to certain non-periodic set ups.
In Section 1.2, we define an extension of the typical two-periodic Aztec diamond, which we refer to as the split two-periodic Aztec diamond. In the split two-periodic model, periodicity is broken along a line which we refer to as the interface of the model. A simulation of the split two-periodic Aztec diamond is shown in Figure 2. We compute the correlation kernel of this new model by extending the methods developed in [BD19]. The statement of the correlation kernel is found in Section 2.1. The rest of our discussion in Section 2 relates to defining the macroscopic regions and computing the local asymptotic of the model away from the boundaries and interface.
Figure 2. Simulated split Aztec diamond where $N=200$, $\alpha=1/4$, and $\beta=1/2$. The tiles are colored by an 8 color gray scale to accentuate the smooth region. Original code for simulation was provided by Sunil Chhita. Simulations are sampled exactly from the distribution using domino-shuffling, see [Pro03].
@@ 1.2. The Split Two-Periodic Aztec Diamond
and
(2) $\mathtt{B}^{\text{Az}}_{n}=\Big\{(2j,2k+1):0\leq j\leq n\text{ and }0\leq k\leq n-1\Big\}.$
The Aztec diamond is a bipartite graph, so we write the set of all vertices as $\mathtt{V}_{n}^{\text{Az}}=\mathtt{W}^{\text{Az}}_{n}\sqcup\mathtt{B}^{\text{Az}}_{n}$. The edges are given by,
(3) $\mathtt{E}^{\text{Az}}_{n}=\Big\{((2j+1,2k),(2j+1\pm 1,2k+1)):0\leq j\leq n-1,0\leq k\leq n-1\Big\}\\ \cup\Big\{((2j+1,2k),(2j+1\pm 1,2k-1)):0\leq j\leq n-1,1\leq k\leq n\Big\}.$
00111122223333444455556666
0011112222333344445555666677778888
@@
where $0<\alpha,\beta<1$ are constants. The last case in the above formula indicates the the weight is 0 if the vertices are not neighbors. If $\alpha=\beta$, this model recovers the usual two-periodic Aztec diamond. If $\alpha=\beta=1$ we recover the uniform Aztec diamond, however we will generally assume that the constants are not equal to 1. The split two-periodic Aztec diamond of size $2N=4$ is shown in Figure 4. We will refer to the line were the weighting changes as the interface of the model. In this coordinate convention, the interface occurs along the vertical line $x=2N$.
[…]
@@
Figure 4. The Aztec diamond of size $2N=4$ with coordinates given by equations (1)-(2) and weights given by equation (4). All unlabeled edged have weight 1.
1.3. A Related Paths Process and the Correlation Kernel
To study the split two-periodic Aztec diamond we will extend the work of Berggren and Duits in [BD19]. Their work focuses on a class of non-intersecting paths models which, under appropriate conditions on the edge weights, are equivalent to certain Aztec diamond dimer models. The equivalence of the Aztec diamond to certain non-intersecting paths models was initially realized by Johansson [Jo03]. Here we briefly introduce the non-intersection paths process which is equivalent to the split two-periodic Aztec diamond. Section 5 details the equivalence between these processes.
We will consider a graph which we will refer to as the BD-paths graph of size $n=2N$. The graph has the following vertex set,
$\mathtt{V}^{\text{BD}}_{n}=\Big\{(j,k):0\leq j\leq 2n,\,-n-1\leq k\leq-1\Big\}$
and the following edge set,
$\mathtt{E}^{\text{BD}}_{n}=\Big\{((j,k),(j+1,k)):0\leq j<2n,\,-n-1\leq k\leq-1\Big\}\\ \cup\Big\{((2j,k),(2j,k-1)):1\leq j\leq n,\,-n\leq k\leq-1\Big\}\\ \cup\Big\{((2j,k),(2j+1,k-1)):0\leq j\leq n,\,-n\leq k\leq-1\Big\}.$
The edges are split up into horizontal, vertical, and diagonal edges. The BD-paths graph of size $4$ is show in Figure 5. On the graph we assume that there are $n$ non-intersecting paths that start at the points (0,k)(0,-k) for $n\leq k\leq-1$ and that end at the points (2j,n1)(2j,-n-1) for $1\leq j\leq 2n$. We say that there is a point at the vertex $(j,k)$ if a path crosses through the vertex and the point $(j,k+1)$ does not lie on the same path. An example configuration of the non-intersecting paths process is show in Figure 6. Figure 6 also shows the equivalent Aztec diamond dimer covering (more details on this can be found in Section 5).
[…]
Figure 5. An example of the BD-paths graph for $n=4$. The graph is also equipped with the edge weights that make it equivalent to the split two-periodic Aztec diamond, see Section 5 for more details. Unlabeled edges have weight 1.
Figure 6. An example of the non-intersecting path process for $n=4$. The equivalent dimer covering of the size 4 Aztec diamond is also pictured.
To make the paths process have the same probability measure as the split two-periodic Aztec diamond we assign the following weights to the edges: to the horizontal edges,
$\text{wt}((j,k),(j+1,k))=1$
to the vertical edges,
[…]
and to the diagonal edges,
[…]
As stated in [BD19], this paths process of size $n=2N$ is a determinantal point process (as is the Aztec diamond dimer model), meaning there exists a matrix $\mathbb{K}_{N}$ such that
$\mathbb{P}\big(\text{points at }(m_{1},u_{1}),\,\dots,\,(m_{k},u_{k})\big)=\det\big(\mathbb{K}_{N}(m_{i},u_{i};m_{j},u_{j})\big)_{i,j=1}^{k}.$
We call this matrix the correlation kernel of the process.
Remark 1.1.
The correlation kernel of a determinantal point process is, in general, not unique. For those familiar with the Kasteleyn approach to dimer models (see [Ken2009lectures] and [Gor20] for overviews), the inverse Kasteleyn is a correlation kernel of the dimer model, however the Kasteleyn matrix depends on a choice of Kasteleyn weighting (and is thus not unique, though the matrices are relatable).
Remark 1.2.
We will use the coordinate notation from the paths process when defining the correlation kernel of the split two-periodic Aztec diamond in Section 2.1. One can, heuristically, think of these coordinate changes as scalar transformations, even though the bijection between the models is a bit more complicated (see Section 5). One can also refer to the work in [CD23] for a detailed explanation on the relationship between these processes and their kernels.
@@ 2. Summary of Results
In this section, we will detail the main results of this work. This starts with the statement of the correlation kernel of the split two-periodic Aztec diamond, which we later compute by extending the methods of Berggren and Duits [BD19]. We then relate this kernel to the kernel of the typical two-periodic Aztec diamond model. Lastly, we classify the macroscopic regions of the model by way of computing the local asymptotics. In particular, we emphasize the differences in the local asymptotics of the split two-periodic model compared to the typical two-periodic model.
@@ 2.1. An Integral Representation of the Correlation Kernel
$\phi_{\varepsilon,1}(z)$ $=\begin{pmatrix}1&\varepsilon^{2}z^{-1}\\ \varepsilon^{-2}&1\end{pmatrix}$, $\phi_{\varepsilon,2}(z)$ $=\frac{1}{1-z^{-1}}\begin{pmatrix}1&\varepsilon^{2}z^{-1}\\ \varepsilon^{-2}&1\end{pmatrix}$,
$\phi_{3}(z)$ $=\begin{pmatrix}1&z^{-1}\\ 1&1\end{pmatrix}$, $\phi_{4}(z)$ $=\frac{1}{1-z^{-1}}\begin{pmatrix}1&z^{-1}\\ 1&1\end{pmatrix}$,
@@
(5) $\Phi_{\varepsilon}(z)=\phi_{\varepsilon,1}(z)\phi_{\varepsilon,2}(z)\phi_{3}(z)\phi_{4}(z).$
@@
(6) $\Phi_{\varepsilon}(z)=E_{\varepsilon}(z)\begin{pmatrix}r_{\varepsilon,1}(z)&0\\ 0&r_{\varepsilon,2}(z)\end{pmatrix}E_{\varepsilon}(z)^{-1}.$
@@
(8) $r_{\varepsilon,2}(z)=\frac{1}{(z-1)^{2}}\left((z+1)^{2}+2z(\varepsilon^{2}+\varepsilon^{-2})-2(\varepsilon+\varepsilon^{-1})\sqrt{z^{3}+(\varepsilon^{2}+\varepsilon^{-2})z^{2}+z}\right).$
We chose the branch cuts for the above to be $(-\infty,-\varepsilon^{-2}]\cup[-\varepsilon^{2},0]$ and for $z>0$ we choose the positive square root. We also adopt the notation,
@@
$=\begin{pmatrix}\frac{1}{2}+\frac{z(\varepsilon^{2}-1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}&\frac{\varepsilon^{2}(z+1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}\\ \frac{z(z+1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}&\frac{1}{2}-\frac{z(\varepsilon^{2}-1)}{2\sqrt{z(z+\varepsilon^{2})(1+\varepsilon^{2}z)}}\end{pmatrix}$.
which uses the same branch cuts as above. Lastly we define the function,
(11) $g_{\alpha,\beta}(z)=\frac{2z(1+\alpha^{2}\beta^{2})+\beta^{2}(z^{2}+1)+\alpha^{2}(z^{2}+1)}{4\sqrt{(z+\alpha^{2})(1+\alpha^{2}z)(z+\beta^{2})(1+\beta^{2}z)}}.$
Note that $g_{\alpha,\beta}(z)=g_{\beta,\alpha}(z)$, and $g_{\alpha,\alpha}(z)=\nicefrac{{1}}{{2}}$. Assuming $\beta>\alpha$, $g_{\alpha,\beta}(z)$ has branch cuts $\alpha^{-2},-\beta^{-2}]\cup[-\beta^{2},-\alpha^{2}]$. We are now prepared to state the main theorem of this paper,
@@
Let $N\leq\xi,\xi^{\prime}\leq-1$ and $0<m<N$. The split two-periodic Aztec diamond of size $2N$ has correlation kernel given by
(12) $\Big[\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\alpha}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
(13) $\Big[\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(w)F_{\alpha,1}(w)F_{\beta,2}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
when $\nicefrac{{N}}{{2}}<m^{\prime}<N$. For both cases, if $m\leq\nicefrac{{N}}{{2}}$ then $\varepsilon=\alpha$ and if $m>\nicefrac{{N}}{{2}}$ then $\varepsilon=\beta$. Additionally, $\gamma_{1}$ is a contour surrounding $1$ and not 0, while $\gamma_{0,1}$ is a contour surrounding 0 and $\gamma_{1}$. Both are positively oriented. We use the index $j$ to denote the matrix row and the index $i$ to denote the matrix column.
[…]